P4- 31 (problem)

Financial statement analysis

The annual sales for Salco, Inc. were $4.51 million last year. The firms end of year balance sheet was as follows (will post) Salco's income statement for the year was as follows (will post)

A. Calculate Salco's total asset turnover, operating profit margin, and operating return on assets.

B. Salco plans to renovate one of it's plants and the renovation will require an added investment in plant and equipment of $1.01 million. The firm will maintain it's present debt ratio of 50 percent when financing the new investment and expects sales to remain constant. The operating profit margin will rise 13.8 percent. What will be the new operating return on assets ratio (ie; net operating income/ total assets) for Salco after the plants renovations?

C. Given that the plant renovation in part (b) occurs and Salco's interest expense rises by $54,000 per year, what will be the retun earned on the common stockholders investment? Compare this rate of retun with that earned before the renovation. Based on this compairson, did the renovation have a favorable effect on the profitability of the firm?

Anwer: (Round to two decimal places)

Income Statement

Sales    $4,510,000
Less: Cost of goods sold    (3,490,000)
Gross profit    $1,020,000
Less: Operating expenses    (493,000)
Net operating income    $527,000
Less: Interest expense    (99,000)
Earnings before taxes    $428,000
Less: Taxes (35%)    (149,800)
Net income    $278,200

End of Year Balance Sheet

Current assets $503,000 Liabilities $999,500
Net fixed assets 1496000 Owners' equity 999500
Total Assets $1,999,000 Total

$1,999,000
Answer #1
A. Total asset turnover= Sales/Total assets
ie. 4510000/1999000=
2.26
times
Operating profit margin= Operating profit(ie. EBIT)/Sales
ie. 527000/4510000=
11.69%
Operating return on assets ratio=
Net operating income(EBIT)/ total assets)
527000/1999000=
26.36%
ie.Opg.ROA=Total asset Turnover*Opg.PM
ie.2.26*11.69%= 26.36%
b.New Total asset turnover=
ie 4510000/(1999000+1010000)=
1.50
times
So,
Opg.ROA= New Total asset Turnover*Increased Opg.PM
ie. 1.5*13.8%=
20.7%
C. Net income=((Sales*Increased Opg.profit margin)-Interest expense)*(1-Tax rate)=
((4510000*13.8%)-(99000+54000))*(1-35%)=
305097
Return on equity=(revised)Net income/New Owners' equity
ie.305097/1504500=
20.28%
(for new equity , see balance sheet after renovation)
Before renovation
Return on equity=Net income/Owners' equity
ie.278200/999500=
27.83%
Balance Sheet
Before renovation
Current assets $503,000 Liabilities $999,500
Net fixed assets 1496000 Owners' equity 999500
Total Assets $1,999,000 Total $1,999,000
Given that
the firm will maintain it's present debt ratio of 50 percent when financing the new investment
After renovation
Liabilities $1,504,500
Owners' equity $1,504,500
New Assets total $3,009,000 Total $3,009,000
Ratios Before renovation After renovation Comments
TATO 2.26 1.5 Decreased
Opg.PM 11.69% 13.80% Increased
Opg.ROA 26.36% 20.70% Decreased
ROE 27.83 20.28% Decreased
Net Income margin 6.17% 6.76% Increased
Even though profit margins increased,both return on assets and return on equity decreased, indicating declining asset utilisation , ie. $ sales generated per $ of asset used has decreased--as sales remain constant ,despite increase in assets.Opreational efficiency has increased, as seen by the increased profit %
So, it can be concluded that
the renovation has a favorable effect on the profitability of the firm
but the asset utilisation needs to improve.